x/2 + y/3=7/6
2x/3+3y/4=5/4
Is there something I can do here so I can get rid of the fractions?

Answer by rothauserc(4718)

(1) x/2 +y/3 = 7/6
:
(2) 2x/3 +3y/4 = 5/4
:
In each equation, multiply both sides of the = by the least common multiple(lcm)
:
For equation (1), the lcm is 6, therefore equation 1 becomes
:
3x +2y = 7
:
Foe equation (2), the lcm is 12, therefore equation 2 becomes
:
8x +9y = 15
: